OpenStudy (brittany):

i was saloving for a quadratice equation and i got down to the numerator equal to -b+-8 and the denominator as 2, how do i simplify

OpenStudy (anonymous):

you have to leave it as it is

OpenStudy (anonymous):

do you have a value for b?

OpenStudy (amistre64):

in order to help we would do better to see your original equation

OpenStudy (brittany):

b=2

OpenStudy (anonymous):

ok so the answer is either -5 or 3

OpenStudy (brittany):

x^2+2x-15=0

OpenStudy (anonymous):

so the coefficients are 1, 2 and -15

OpenStudy (brittany):

yes

OpenStudy (amistre64):

now lets see if your work thru is good: \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\frac{-2\pm\sqrt{2^2-4(1)(-15)}}{2(1)}\] \[\frac{-2\pm\sqrt{4+60}}{2}\] \[\frac{-2\pm\sqrt{64}}{2}\] \[\frac{-2\pm8}{2}\]

OpenStudy (brittany):

yes :)

OpenStudy (amistre64):

so you did fine except for you didnt plug in your "2" into that top "b" :)

OpenStudy (brittany):

yes now how do i finish tht problem? divide -1 and 8 by two?

OpenStudy (brittany):

-2 ans 8 sorry

OpenStudy (amistre64):

the +- sign means we have 2 equations, its just a compact way to write this: \[\frac{-2+8}{2}\text{ and }\frac{-2-8}{2}\] then you can simplify

OpenStudy (anonymous):

what is the question exactly?

OpenStudy (anonymous):

there are two solutions to the equation

OpenStudy (amistre64):

-2+8 = 6 ; 6/2 = 3 -2-8 = -10; -10/2 = -5

OpenStudy (brittany):

thankyou :)

OpenStudy (anonymous):

no problem

OpenStudy (amistre64):

i find that when I am learning new things, that I have to remember not to forget the old stuff I learned ;)

OpenStudy (anonymous):

"remember not to forget" = lol